OpenStudy (anonymous):

express each angle measure degrees: 4.07rad

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

really? -_- It's like you don't even learn Jake XD Just multiply the radian measure by \(\LARGE \frac{180^o}{\pi}\approx57.2957795131\) and you'll get the measure in degrees.

OpenStudy (anonymous):

ahaahahah, sorry,, wrong given,, that was supposed to be for another question,,

terenzreignz (terenzreignz):

Jake talaga oh

OpenStudy (anonymous):

\[\frac{ 4\pi }{ 3 }\]

terenzreignz (terenzreignz):

It's the same bloody thing, except there's no need for that approximation. Just multiply by \(\LARGE \frac{180^o}\pi\). If you don't like it, just replace the \(\pi\) with 180, and then simplify.

OpenStudy (anonymous):

should i just reciprocal 180/pi?

OpenStudy (anonymous):

damn,, that was convert to mins

terenzreignz (terenzreignz):

radians--> degrees, multiply \(\LARGE \frac{180^o}\pi\) degrees--> radians, multiply \(\LARGE \frac\pi{180^o}\)

terenzreignz (terenzreignz):

convert to minutes? D:

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

To degrees first, what do you get?

OpenStudy (anonymous):

uhm, pellet, someone's calling me be right back

OpenStudy (anonymous):

pelleeeeeeeeeeeeeeeeeeeeeeets!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

pelleeeeeeeeeeeeeeets and stuff

terenzreignz (terenzreignz):

ehh?

terenzreignz (terenzreignz):

Convert the angle to degrees first.

OpenStudy (anonymous):

i already got it,, no need to solve, :3 anyways i need to save time,

terenzreignz (terenzreignz):

bleargh

OpenStudy (anonymous):

how can i convert rad to degrees again,, i

terenzreignz (terenzreignz):

read. up.

OpenStudy (anonymous):

example is 4.07 rad

terenzreignz (terenzreignz):

bloody hell, that's on the top of this post :P

terenzreignz (terenzreignz):

It's like the first thing I said here XD

terenzreignz (terenzreignz):

Do you even listen *Jake*? XD

OpenStudy (anonymous):

uhm,, in the equation 180/pi,, is there rad next to pi?

terenzreignz (terenzreignz):

sure, why not.

OpenStudy (anonymous):

oh,,why is there rad next to pi?

terenzreignz (terenzreignz):

there doesn't really have to be, radians are pretty much real numbers.

terenzreignz (terenzreignz):

unlike degrees, which are strictly angle measures.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i get 233.2

terenzreignz (terenzreignz):

acceptable.